The amount of money invested in a certain account increases according to the following function, where
y0
is the initial amount of the investment, and
y
is the amount present at time
t
(in years).
=yy0e0.045t
After how many years will the initial investment be doubled? Do not round any intermediate computations, and round your answer to the nearest tenth.

Answer by Boreal(15235)

this is the continuous compounding formula
y=yo*e^(0.045t), y/yo=e^(0.45t)
when it doubles, y/yo=2
so 2=e^(0.045t)
ln both sides
ln2=0.045t
t=15.4 years (rule of 70, where interest rate (in %)* time(years)=70 as a quick way to get a sense of doubling time.